Field Notes
<--2017-11-10 14:45:00 | 2017-11-30 10:20:00-->Other sites visited today: Sherman Wetland | Bouldin Corn | East End | Twitchell Alfalfa
Phenocam link: Searching...
Mayberry: 2017-11-17 10:10:00 (DOY 321)
Author: Daphne Szutu
Others: Joe
Summary: Regular data collection/cleaning, collected water samples for Isabel, re-wired SWin to new mux channels, cleaned fluorometer
2017-11-17 Mayberry We arrived at 10:10. It was sunny and clear, a beautiful day in the Delta. There were some low clouds on the horizon. The cattails are chewed but still green. The solar charge controller had ants, so we put in some mothballs. The site has started losing power some nights, so we should swap the solar panels or batteries. Joe cleaned the fluorometer and emptied out the bubbles after re-deployment. Joe collected water samples near the fluorometer for Isabel at 10:45. He was already in the water so it was a bit mucky. SWin data has been spiking. I moved SWin from channel 13 to 29 and updated the program, MB_Tule_20171117.CR1. I downloaded USB, met, and camera data. When the Licor time was 10:34, I changed it to 10:39. We cleaned the EC/rad sensors and the outside of the camera enclosure. Wash reservoir was mostly full, and we topped it off. LI-7500 readings: CO2 411.4 ppm H2O 443.5 mmol/m3 T 18.4 C P 102.1 kPa SS 103.8 --> 104 after cleaning LI7700 readings: CH4 1.95 ppm RSSI 59.1 --> 75 after cleaning The medium waders have leaks in the right knee and butt and need to be patched. We left at 10:55. |
1 photos found
20171117MB_fluorometer.JPG ( 2017-11-17 10:24:19 ) Full size: 1080x1440
Fluorometer before cleaning
11 sets found
7 MB_flux graphs found
No data found for MB_met.
28 MB_tule graphs found
Can't check MBWPEESWGT_mixed
1 MB_cam graphs found
Can't check MBEEWPSWEPGT_cam
Can't check SIMBTWEE_met
1 MB_minidot graphs found
Can't check MBSWEEWPBCBA
1 MB_processed graphs found
Can't check MB_fd
1 MB_fd graphs found